OpenStudy (anonymous):

find the three consecutive even integers such that the first times the third is 2 greater than 5 times the second?

OpenStudy (ganpat):

let three consecutive even integers be, x, x+2 and x+4 Therefore, x(x+4) = 5(x+2) +2 x2 + 4x = 5x + 10 +2 x2 - x -12 = 0 x2 -4x + 3x -12 = 0 x (x -4) + 3 (x - 4) = 0 (x-4) (x+3) =0 x= -3 and x =4...